OpenStudy (anonymous):

how would i write,three more than seven times a number is that number increased by 38, i have to write an equation to represent the problem

OpenStudy (anonymous):

3+7x=x+38

OpenStudy (anonymous):

how can i figuire it out

OpenStudy (anonymous):

remember that is and = mean the same thing

OpenStudy (anonymous):

the perimeter of a retangle room is 166 feet. find the length and width of the room if the length is 5 feet longer than twice the width

OpenStudy (anonymous):

You have to remember the formula for perimeter of a rectangle to figure this out

OpenStudy (anonymous):

let the width W = w ft then length L = (w + 5) ft Perimeter of rectangle = 2L + 2W ATQ, 166 = 2(w + 5) + 2(w) 166 = 2w + 10 + 2w 166 - 10 = 4w 156/4 = w 39 = w so width = 39 ft length = (39 + 5) = 44 ft